# Leveraging Large Language Models for Collective Decision-Making

### v2

@@ Abstract.
In various work contexts, such as meeting scheduling, collaborating, and project planning, collective decision-making is essential but often challenging due to diverse individual preferences, varying work focuses, and power dynamics among members. To address this, we propose a system leveraging Large Language Models (LLMs) to facilitate group decision-making by managing conversations and balancing preferences among individuals. Our system aims to extract individual preferences from conversations and suggest options that satisfy the preferences of the members. We specifically apply this system to corporate meeting scheduling. We create synthetic employee profiles and simulate conversations at scale, leveraging LLMs to evaluate the system performance as a novel approach of conducting a user study. Our results indicate efficient coordination with reduced interactions between the members and the LLM-based system. The system refines and improves its proposed options over time, ensuring that many of the members’ individual preferences are satisfied in an equitable way. Finally, we conduct a survey study involving human participants to assess our system’s ability to aggregate preferences and reasoning about them. Our findings show that the system exhibits strong performance in both dimensions.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; Under Review; price: 15.00isbn: 978-1-4503-XXXX-X/18/06conference: The 27th ACM SIGCHI Conference on Computer-Supported Cooperative Work & Social Computing; November 9-13, 2024; San Jose, Costa Ricaccs: Human-centered computing Collaborative and social computingccs: Computing methodologies Cooperation and coordinationccs: Computing methodologies Multi-agent systemsccs: Information systems Social networks
@@ 1. Introduction
In particular, in work environments such as organizing meetings, collective decision-making tasks have been a productivity bottleneck (Saha and Iqbal, 2023) due to their inherent diversity of preferences across members from possibly different organizations (Crawford and Veloso, 2007; Crawford, 2009; Ephrati et al., 1994; BenHassine and Ho, 2007; Brzozowski et al., 2006; Mok et al., 2023), and have become more complex in the hybrid work setting(Yang et al., 2022b, 2023, a). Besides traditional methods using calendars (e.g., Doodle, when2meet, MS Outlook, etc.), copilots such as Calendar.help (Cranshaw et al., 2017) have been proposed to aid with meeting scheduling and distributed decision-making (e.g., (Yardi et al., 2005; Álvarez Márquez and Ziegler, 2016; Cao et al., 2018)) across individuals within or across organizations. However, these approaches have been tailored to specific sub-problems and cannot encompass the open-ended preferences and constraints the members may have. [^1] See Section 9 for a thorough comparison with these related works.
@@
In this paper, we leverage the power of LLMs to create a system that assists collective decision-making within a group, such as selecting a meeting time or deciding a venue for a group event. The system’s main goal is to guide the group discussion towards an option that respects the opinions of the members. LLMs play a crucial role in this system by extracting individual preferences from the members, consolidating these preferences, and subsequently proposing options. Importantly, the system also utilizes the LLMs to provide the reasoning behind each suggestion, ensuring transparency and understanding among the group members.
Specifically, our system is designed with four key components: (i) a knowledge graph, which serves as a repository for background information such as metadata and a social graph; (ii) an intent extraction module, tasked with discerning individual preferences from the dialogues between the system and the members; (iii) a coordination module, responsible for generating options that align with the members’ preferences; (iv) an evaluation module that assesses how well the proposed options meet the members’ preferences. Figure 1 provides a comprehensive overview of our system.
@@
The findings of our large-scale study show a consistent decrease in the number of interactions between the members and the system with each coordination round. The system demonstrates effectiveness in proposing options that increasingly satisfy a greater number of members and fulfill their preferences in a more balanced manner over time. Our survey result indicates that the human participants view the system as having strong capabilities of preference aggregation and reasoning. The participants also expressed their confidence in utilizing LLMs for the coordination of members’ preferences during collective decision-making processes.
@@
Paper Organization. The paper is organized as follows: In Section 2, we perform a requirement analysis for our proposed system, analyzing responses from our participants regarding their characteristics and needs and comparing our system design with existing systems. Section 3 introduces the main components of our system and the collective decision-making process. In Section 4, we focus on group meeting scheduling as a case study for our system. In Sections 5 and 6, we conduct a large-scale simulation study to evaluate our system’s performance, and in Section 7 our study participants evaluate the system’s reasoning abilities. Finally, we discuss our findings in Section 8 and review related literature in Section 9.
Supplemental Information. The Appendix of our paper has been uploaded as supplementary materials. The Appendix details the system implementation prompts, survey design, and experimental results.
@@
Figure 1. Collective decision-making leveraging LLMs. (a) A high-level overview of our proposed system: The LLM-based system processes individual preferences, opinions, and small talk and suggests options that satisfy as many members as possible. (b) A block diagram of our LLM-based system as described in Section 3. The members provide messages $\bm{m}^{t}$ to the intent extraction module, which distills the true preferences $\bm{\theta}^{t}$ from $\bm{m}^{t}$. The coordination module also has access to a knowledge graph $\mathcal{G}$. Subsequently, the coordination module (coordinator) is responsible for suggesting the option set $\mathcal{X}^{t}=\{x_{1}^{t},\dots,x_{K}^{t}\}$, and the evaluation module is responsible for assigning a score $\pi_{j,i}^{t}$ for every member’s preferences $\theta_{i}^{t}$ and every option $x_{j}^{t}$. The decision candidate is selected as the “best” (see Equation 4) option out of the option set $\mathcal{X}^{t}$ and is provided to the coordination module for round $t+1$.
@@ 2. Requirements Analysis
In the section, we aim to understand user characteristics, usage contexts, user needs, and the design of previous systems. We performed a user study in a technology company; we recruited the employees via internal campaigns and mailing lists. The study was approved by the ethics committee within our organization (cf. Section 11), and the participants who completed our study were compensated for 15 US dollars. For more information about the study design, see Appendix D.
@@ 2.1. Participants
We received a total of 45 effective responses. The participants held various roles (e.g., software engineer, product manager, etc.). Most of the participants (78%) had at least 3 years of experience. About half of the participants (55%) indicated that their meetings usually had 3 to 5 attendees, and few of them had meetings with size larger than 10. Regarding the meeting frequency, 28% of the participants had between 1 and 2 meetings per day, 54% of them had between 3 and 5 meetings per day, and the rest had more than 6 meetings per day. Mid-senior employees had the most meetings per day. Figure 2 plots the statistics.
@@ 2.2. Usage Contexts and User Needs
The participants were asked to indicate their intended use of an LLM-based copilot for collective decision-making. The participants could express in free-text but we provided several potential scenarios: (1) meeting scheduling, (2) group discussion for distributed communication application, (3) brainstorming for collaborative applications, and (4) event planning for workplace management. We found that meeting scheduling was considered an important use case by almost all (95%) of the participants, followed by group discussion (94%), brainstorming (78%), and event planning (27%). Two participants also mentioned uses as HR tools and software development. This survey result suggested meeting scheduling as an important scenario of our proposed framework, as well as indicated the essential parameters to study.
@@
2.3. Existing Systems
A closely-related work to ours is Calendar.help (Cranshaw et al., 2017), which designs a meeting scheduling assistant through structured workflows. Specifically, users interact with Calendar.help by email and Calendar.help parses the workflows and schedules meetings by extracting information from the users’ calendars of meetings and tasks where human-in-the-loop interventions may be needed.
@@ 2.3. Existing Systems
Calendar.help consists of microtask execution layers that is based on a combination of machine learning, natural language processing, and heuristic-based components to extract information about the meeting request, details about the attendees, the subject, the time, and the location of the meeting and coordinate the meeting time with the attendees. When the automatic microtasks fail, Calendar.help relies on human-in-the-loop interventions to resolve the scheduling issues.
Our system for collective decision-making shares several features with Calendar.help, but is at the same time significantly different: Specifically, we design four modules to extract individual preferences from dialogues and suggest meeting times, as will be detailed in Section 3. Our system is also dynamic in the sense that it aims to improve its proposed meeting plan through several rounds of conversation with the group members.
@@
On the other hand, Calendar.help supports a limited amount of tasks and uses a complicated architecture, unable to capture diverse open-ended preferences. This work also took significant time (18 months) to complete its user study. By contrast, our work will introduce some innovations: Our model can process open-ended preferences. Its simple architecture of modules that are implemented using prompts allows for quick adaptation to a variety of tasks (e.g., event coordination) beyond meeting scheduling. Moreover, we will adopt an innovative approach of conducting a large-scale user study using LLM-simulated users, therefore reducing costs and facilitating prototyping and assessment of the multi-agent system.
@@ 3. System Design
A leader must make a decision about a topic for a group consisting of $n$ members. We index the members with $i\in[n]\triangleq\{1,...n\}$. Each member $i$ has their individual preferences about the topic. To collect the members’ preferences, the leader organizes a discussion based on which the leader decides for the group. The discussion proceeds at $t=0,1...T$ where $t=0$ is the initialization stage. In each round, the members can share their individual preferences or opinions about the topic and candidate options in their natural languages. A system powered by a large language model (LLM) is utilized to coordinate the discussion, gather opinions, and even suggest decision candidates for the group. We elaborate on the system and the discussion process as follows.
@@ 3.1. Modules
The system consists of four components: (1) an intent extraction module, (2) a coordination module, (3) an evaluation module, and (4) a knowledge graph. Their functions are described below.
  • Intent Extraction Module. The intent extraction module extracts the members’ individual preferences from their free-text messages with the system and outputs them as a list of compact restatements. We write $m_{i}^{t}$ for the conversation between member $i$ and the system at round $t$, which may include briefings or information from previous rounds. We write $\theta_{i}^{t}$ to collect individual preferences extracted from the conversation $m_{i}^{t}$. For simplicity, we use vector notation $\bm{m}^{t}=\left(m_{1}^{t},\dots,m_{n}^{t}\right)$ for the collection of the members’ message at round $t$, and similarly write $\bm{\theta}^{t}=\left(\theta_{1}^{t},\dots,\theta_{n}^{t}\right)$ for their individual preferences.
  • Coordination Module (Coordinator). At each round $t$, the coordination module generates a list of options based on the topic, metadata (e.g., the social relations between the members), and the members’ preferences extracted from their messages at this round (i.e., $\bm{\theta}^{t}$). We denote the set of options at round $t$ as $\mathcal{X}^{t}=\{x_{1}^{t},\dots,x_{K}^{t}\}$, where $K$ represents the number of options that the coordination module produces.
    For exposition purposes, we say that an option $x_{j}^{t}$ satisfies member $i$ ($\theta_{j}^{t}$) when at least one of the preferences of member $i$ is fulfilled by the option $x_{j}^{t}$. We describe what information the coordination module presents for each option later.
  • Evaluation Module. During round $t$, the evaluation module measures the quality of the options $\mathcal{X}^{t}$ proposed by the coordination module based on the degree to which the options incorporate members’ preferences. For every option $x_{j}^{t}\in\mathcal{X}^{t}$ and each member $i$, the module generates a non-negative value $\pi_{j,i}^{t}\geq 0$ that quantifies the degree to which the individual preferences $\theta_{i}^{t}$ of member $i$ are taken into account: The value is increasing in the number of satisfied preferences. We assign $\pi_{j,i}^{t}=0$ exclusively when none of the member $i$’s preferences in $\theta_{i}^{t}$ are fulfilled.
    Based on the values $\{\pi_{j,i}^{t}\}_{i\in[n]}$ for each option $x_{j}^{t}$, we first define the satisfaction ratio of option $x_{j}^{t}$ as the fraction of members for whom this option meets at least one of their individual preferences, which is given by
    (1) $r_{j}^{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\pi_{j,i}^{t}>0\right\}.$
@@
    To investigate whether the members’ preferences are equally considered for each option $x_{j}^{t}$, we define the equity score of option $x_{j}^{t}$, which is given by
@@
    The equity score is set as $1$ if option $x_{j}^{t}$ does not meet any member’s preference, i.e., $r_{j}^{t}=0$. Otherwise, the equity score is the Gini coefficient of the values $\{\pi_{j,i}^{t}\}_{i\in[n]}$ (Gini, 1936). The equity score ranges between 0 and 1.
  • Knowledge Graph. A knowledge graph $\mathcal{G}$ captures background information about the members related to the topic, e.g., their metadata (expertise, responsibilities) and social networks. Incorporating the knowledge graph in the process of coordinating meeting schedules can be vital: It has been shown that these heterogeneities among members have a great impact on meetings in large-scale organizations (Charpignon et al., 2023).

### v3

@@ Abstract.
In various work contexts, such as meeting scheduling, collaborating, and project planning, collective decision-making is essential but often challenging due to diverse individual preferences, varying work focuses, and power dynamics among members. To address this, we propose a system leveraging Large Language Models (LLMs) to facilitate group decision-making by managing conversations and balancing preferences among individuals. Our system aims to extract individual preferences from each member’s conversation with the system and suggest options that satisfy the preferences of the members. We specifically apply this system to corporate meeting scheduling. We create synthetic employee profiles and simulate conversations at scale, leveraging LLMs to evaluate the system performance as a novel approach to conducting a user study. Our results indicate efficient coordination with reduced interactions between the members and the LLM-based system. The system refines and improves its proposed options over time, ensuring that many of the members’ individual preferences are satisfied in an equitable way. Finally, we conduct a survey study involving human participants to assess our system’s ability to aggregate preferences and reasoning about them. Our findings show that the system exhibits strong performance in both dimensions.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; January 1, 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06conference: The 28th ACM SIGCHI Conference on Computer-Supported Cooperative Work & Social Computing; November 9-13, 2024; San Jose, Costa Ricaccs: Human-centered computing Collaborative and social computingccs: Computing methodologies Cooperation and coordinationccs: Computing methodologies Multi-agent systemsccs: Information systems Social networks
@@ 1. Introduction
[…]
@@
In this paper, we leverage the power of LLMs to create a system that assists collective decision-making within a group, such as selecting a meeting time or deciding a venue for a group event. The system’s main goal is to guide the group towards an option that respects the opinions of the members. LLMs play a crucial role in this system by extracting individual preferences from the members, consolidating these preferences, and subsequently proposing options. Importantly, the system also utilizes the LLMs to provide the reasoning behind each suggestion, ensuring transparency and understanding among the group members.
[…]
@@
The findings of our large-scale study show a consistent decrease in the number of interactions between the members and the system with each coordination round. The system demonstrates effectiveness in proposing options that increasingly satisfy a greater number of members and balance their preferences over time. Our survey results indicate that the human participants view the system as having strong capabilities for preference aggregation and reasoning. The participants also expressed their confidence in utilizing LLMs to coordinate members’ preferences during collective decision-making processes.
@@
Figure 1. Collective decision-making leveraging LLMs. (a) A high-level overview of our proposed system: The LLM-based system processes individual preferences, opinions, and small talk and suggests options that satisfy as many members as possible. (b) A block diagram of our LLM-based system as described in Section 3. The members provide messages $\bm{m}^{t}$ to the intent extraction module, which distills the true preferences $\bm{\theta}^{t}$ from $\bm{m}^{t}$. The coordination module also has access to a database $\mathcal{D}$. Subsequently, the coordination module (coordinator) is responsible for suggesting the option set $\mathcal{X}^{t}=\{x_{1}^{t},\dots,x_{K}^{t}\}$, and the evaluation module is responsible for assigning a score $\pi_{j,i}^{t}$ for every member’s preferences $\theta_{i}^{t}$ and every option $x_{j}^{t}$. The decision candidate is selected as the “best” (see Equation 4) option out of the option set $\mathcal{X}^{t}$ and is provided to the coordination module for round $t+1$.
Paper Organization. The paper is organized as follows: In Section 2, we perform a requirement analysis for our proposed system, analyzing responses from our participants regarding their characteristics and needs and comparing our system design with existing systems. Section 3 introduces the main components of our system and the collective decision-making process. In Section 4, we focus on group meeting scheduling as a case study for our system. In Sections 5 and 6, we conduct a large-scale simulation study to evaluate our system’s performance, and in Section 7, our study participants evaluate the system’s reasoning abilities. Finally, we discuss our findings in Section 8 and review related literature in Section 9.
Supplemental Information. The appendix of our paper has been uploaded as supplementary material. The Appendix details the system implementation prompts, survey design, and experimental results.
@@ 2. Requirements Analysis
In this section, we aim to understand user characteristics, usage contexts, user needs, and previous methods or systems. We performed a user study in a technology company; we recruited the employees via internal campaigns and mailing lists. The study was approved by the ethics committee within our organization (cf. Section 11), and the participants who completed our study were compensated with 15 US dollars. For more information about the study design, see Appendix D.
@@ 2.1. Participants
We received a total of 45 effective responses. The participants held various roles (e.g., software engineer, product manager, etc.). Most of the participants (78%) had at least 3 years of experience. About half of the participants (55%) indicated that their meetings usually had 3 to 5 attendees, and few of them had meetings with sizes larger than 10. Regarding the meeting frequency, 28% of the participants had between 1 and 2 meetings per day, 54% of them had between 3 and 5 meetings per day, and the rest had more than 6 meetings per day. Mid-senior employees had the most meetings per day. Figure 2 plots the statistics.
@@ 2.2. Usage Contexts and User Needs
[…]
In an organization, many decisions require input from involved members, who not only indicate their choices but also share the reasons behind them. Our LLM-based system is designed to gather members’ opinions and the rationale behind their choices, ensuring that individual perspectives are considered in group recommendations rather than simply relying on a majority vote.
This is especially important when a group includes participants external to the organization, as coordinators may need extensive communication to understand and respect these participants’ preferences. Our system aims to reduce the communication workload and efficiently identify options that balance the preferences of all members.
The participants were asked to indicate their intended use of an LLM-based copilot for collective decision-making. They were allowed to provide any scenarios in the free text, but we offered several potential cases: (1) meeting scheduling, (2) group discussion for distributed communication applications, (3) brainstorming for collaborative applications, and (4) event planning for workplace management. We found that meeting scheduling was considered an important use case by almost all (95%) of the participants, followed by group discussion (94%), brainstorming (78%), and event planning (27%). Two participants also mentioned the use of HR tools and software development. This survey result suggested meeting scheduling as an important scenario of our proposed framework, and we will use it for our case study.
[…]
Survey participants, who serve as administrators and frequently coordinate numerous events, shared their coordination experiences in detail. They primarily organize frequent meetings, team events, and executive gatherings, dedicating around 70% of their weekly work time to these tasks. These events typically involve 2 to 20 participants, with about 10% from external organizations. Their common scheduling challenges are time zone differences and often extensive communication. The administrators also indicate several decision-making factors, such as executive attendance, travel, budget, and space availability, balanced by prioritizing meeting owners’ needs.
During their coordination process, commonly used tools include email clients, team messaging applications, note-taking applications, and office suite tools (such as spreadsheets) valued for their integrated functionality. On average, the administrators propose three scheduling options. Although 25% of plans may not work for everyone, attendance remains high at 90%, and the administrators express their confidence that preferences are considered equitably. The administrators highlight additional challenges, including the need for supportive scheduling tools, educating others on scheduling protocols, and accommodating executives’ requests, emphasizing the demand for greater flexibility. This underscores a clear use case for our system, which aims to address these specific coordination and scheduling needs effectively.
@@
2.3. Existing Methods or Systems
[…]
The surveyed administrators highlighted commonly used coordination tools, including email clients (Outlook, Gmail, Thunderbird, etc.), instant messaging apps (Teams, Slack, Google Chat, etc.), note-taking tools (Loop, Google Notes, etc.), and spreadsheet tools (Excel, Google Sheets, etc.). Additional tools like Doodle, When2Meet, and Outlook facilitate the collection of time slot preferences and help identify a majority, focusing primarily on availability statistics with minimal emphasis on participants’ reasons for their choices. By contrast, our system is designed to offer greater interpretability by integrating individual preferences and rationales into its recommendations. Given this focus, a direct comparison between our proposed system and these popular polling methods is less meaningful, as our system processes the nuanced preferences behind choices rather than treating each participant’s selection with equal weight.
Additionally, these traditional tools are less effective when some participants are external to the organization, as coordinators often cannot directly access external participants’ calendars. This limitation necessitates extensive back-and-forth communication (e.g., emails), which can cause delays. Our system addresses this challenge by streamlining the process of collecting and aggregating participants’ preferences, reducing communication overhead.
@@ 2.3. Existing Methods or Systems
A closely related work to ours is Calendar.help (Cranshaw et al., 2017), which designs a meeting scheduling assistant through structured workflows. Specifically, users interact with Calendar.help by email and Calendar.help parses the workflows and schedules meetings by extracting information from the users’ calendars of meetings and tasks where human-in-the-loop interventions may be needed. [^2] For more details, Calendar.help consists of microtask execution layers that are based on a combination of machine learning, natural language processing, and heuristic-based components to extract information about the meeting request, details about the attendees, the subject, the time, and the location of the meeting and coordinate the meeting time with the attendees. When the automatic microtasks fail, Calendar.help relies on human-in-the-loop interventions to resolve the scheduling issues.
@@
Our system shares several features with Calendar.help, but is also significantly different: Specifically, we design four modules to extract individual preferences from dialogues and suggest meeting times, as will be detailed in Section 3. Our system is also dynamic in the sense that it aims to improve its proposed meeting plan through several rounds of conversation with the group members. On the other hand, Calendar.help supports a limited amount of tasks and uses a complicated architecture, unable to capture diverse free-text preferences. This work also took significant time (18 months) to complete its user study. By contrast, our work will introduce some innovations: Our model can process free-text preferences. Its simple architecture of modules that are implemented using prompts allows for quick adaptation to a variety of tasks (e.g., event coordination) beyond meeting scheduling. Moreover, we will adopt an innovative approach of conducting a large-scale user study using LLM-simulated users, therefore reducing costs and facilitating prototyping and assessment of the multi-agent system.
[…]
Another related work is the interactive LLM-based decision support system developed in (Lawless et al., 2024), which also focuses on user preferences, using meeting scheduling as a case study. In their system, the role of LLMs is to translate users’ scheduling constraints into code, which is then input into a scheduling optimization package to find feasible meeting times. In contrast, our system uses LLMs to directly process members’ preferences and generate suggestions. We consider our approach complementary to the system in (Lawless et al., 2024). Specifically, when constraints involve complex relationships and logical manipulations, the system in (Lawless et al., 2024) can leverage optimization packages to effectively and precisely account for user preferences. However, this approach does not provide explanations for the suggestions, as optimization packages typically output only results without rationale. Moreover, optimization packages may not always be available for a given collective decision problem; our system avoids this limitation by using LLMs to directly aggregate preferences.
@@ 3. System Design
[…]
A group of $n$ members are to make a decision; for example, an engineering team wants to schedule a meeting with external customers to finalize their product solution. We index the members with $i\in[n]\triangleq\{1,...n\}$. Each member $i$ has their individual preferences about the decision. A system powered by a large language model (LLM) is utilized to assist the group in decision-making by collecting and understanding the members’ preferences and accordingly suggesting options for the group. This coordination process proceeds at $t=0,1...T$ where $t=0$ is the initialization before the coordination process. At each round, the system shows each member the current decision of the group, and each member can privately communicate with the system about their individual preferences or opinions in their natural languages. We provide the specifics of the system and the process as follows.
@@ 3.1. Modules
[…]
    For simplicity, we use vector notation $\bm{m}^{t}=\left(m_{1}^{t},\dots,m_{n}^{t}\right)$ for the collection of all the members’ messages at round $t$, and similarly write $\bm{\theta}^{t}=\left(\theta_{1}^{t},\dots,\theta_{n}^{t}\right)$ for their individual preferences.
[…]
    To understand how the members’ preferences are served by the system, we define three metrics that characterize the distribution of the values $\{\pi_{j,i}^{t}\}_{i\in[n]}$ among the members for any option $x_{j}^{t}$. We first define the satisfaction ratio of option $x_{j}^{t}$ as the fraction of members for whom this option meets at least one of their individual preferences, which is given by
    (1) $r_{j}^{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\pi_{j,i}^{t}>0\right\},$
    which lies between 00 and $1$. For expositional purposes, we say that an option $x_{j}^{t}$ satisfies member $i$ when at least one of their individual preferences ($\theta_{j}^{t}$) is fulfilled by the option, i.e., $\pi_{j,i}^{t}>0$.
@@
    To investigate whether the members’ preferences are considered in a balanced manner, we define the equity score of option $x_{j}^{t}$, which is given by
@@
    The equity score is set as $1$ if option $x_{j}^{t}$ does not meet any member’s preference, i.e., when $r_{j}^{t}=0$. Otherwise, the equity score is the Gini coefficient of the values $\{\pi_{j,i}^{t}\}_{i\in[n]}$ (Gini, 1936). The equity score is between 0 and 1.
[…]